OpenStudy (dmezzullo):

Solve for x: 1 over 4 (2x − 14) = 4 x = 15 x = 15 over 2 x = 9 x = 11

OpenStudy (dmezzullo):

wrong one

OpenStudy (anonymous):

1=16(2x-14) 1=32x-224 1+224=32x 225/32=x